OpenStudy (anonymous):

Solve using square roots. x^2 - 32 = 17

OpenStudy (anonymous):

x^2 - 32 + 32 = 17 +32 x^2 = 49 sqrt(x^2) = sqrt(49) x = 7

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

Your welcome! :)